Find the area of a circle whose circumference is 18pi diameter.
Answer by vidyamanohar(13)

circumference=pi x diameter =18pi
so, diameter = 18
the radius is half the diameter, so, radius =9
Area of the circle= pi(radius)^2=pi(9)^2=81 pi